# BTGenBot: Behavior Tree Generation for Robotic Tasks with Lightweight LLMs

### v1

@@ Abstract
Index Terms:
robotics, robot behavior, behavior tree, large language models
@@ I Introduction
An example of these representations is behavior trees (BTs). Behavior trees have gained significant traction in the field of robotics, offering a structured and scalable approach to managing complex robot behaviors. Originating from the video game industry, BTs have found relevance in robotics due to their ability to handle high-level decision-making and control in dynamic environments. A key player in this transition is the BehaviorTree.CPP library [^1] https://www.behaviortree.dev/. This library has become the de facto standard in the Robot Operating System 2 (ROS2) ecosystem, largely due to its inclusion in Navigation2 [12]. This integration has facilitated the development of sophisticated navigation behaviors, contributing to the broader adoption and standardization of BTs in robotics software design.
@@
Recently, large language models (LLM) trained on large amounts of data have emerged as a promising tool to manage this challenge and deal with the complex knowledge required for task planning. These models have demonstrated remarkable success in robot planning from a set of natural language instructions. Additionally, an LLM can represent commonsense priors (e.g., visit the colder locations first) and comprehend spatial relationships (e.g., move behind the sofa). However, such large models present a significant limitation: their substantial hardware resource requirements prevent their direct deployment on robots.
In this work, we explore an alternative approach, focusing on compact large language models, specifically those with 7 billion parameters. We posit that these smaller models, while less resource-intensive, can generate behavior trees for robots with comparable effectiveness to their larger counterparts. Our main contributions are: (i) a dataset of 600 behavior trees paired with a natural language description of their tasks, (ii) a system to evaluate and validate the performance of the generated BT over nine different tasks including navigation and manipulation, and (iii) a comparative analysis of three different LLM and their fine-tuned versions. All the material used in this work, including the dataset, the complete prompts of both evaluation phases, and the source code of the validator are available at https://github.com/AIRLab-POLIMI/BTGenBot.
@@ II Related work
A pre-trained foundation model (PFM) refers to a complex neural network-based model trained on extensive open-source datasets, boasting billions of parameters [23]. This robust architecture serves as a versatile backbone that can be further fine-tuned for diverse tasks across various domains. Recently, several PFMs specifically trained on massive text datasets (i.e., large language models) have been proposed such as the most recent one GPT-4 [1] by OpenAI or LLaMA by Meta AI [18]. In this work, we rely on LLaMA released by Meta AI, a family of foundation models trained on an open-source dataset with trillions of tokens. In particular, we employed LLaMA-2 [19], trained on 2 trillion tokens, and with a context length of 4000 tokens, double the length of its predecessor. Three model sizes are available: 7, 13, and 70 billion parameters.
@@
Behavior tree generation. An example of behavior tree generation using LLM is the work done in [11]. In this paper, the authors exploit a transformer-based LLM fine-tuned from the Stanford Alpaca 7B model [17] to generate behavior trees from a text description. They further fine-tuned this model with natural language descriptions generated using text-davinci-003 from behavior trees created in the previous step. In summary, they use two different LLMs to generate both the behavior trees and the task description. While interesting in proving that LLMs can be used to generate BTs, this approach has several limitations. Fine-tuning an LLM with a dataset generated by the LLM itself causes a data bias only partially mitigated by the use of an auxiliary LLM to generate the descriptions. Moreover, self-generated data can be noisy and, in this configuration, errors propagate from the data generation phase to the training phase. Differently from the solution proposed in [11], we create our instruction-following dataset by collecting behavior trees from open-source robotics projects. This allows us to fine-tune our model with BTs already tested and effectively used in real projects. Moreover, we use GPT-3.5 only to generate descriptions for each corresponding behavior tree.
@@
Finally, concerning the validation of the above approach, the only metric used is the visual evaluation by a group of human experts. The evaluators were presented with a pair of behavior trees and had to indicate which one was generated. This evaluation, while interesting and useful to asses the general capabilities of the model, is rather limited. This method does not consider the correctness from a syntactical and semantic point of view, or any performance metric, such as the generation time and the hardware resources used. In contrast, we tested our system with a collection of metrics described in Section III.
@@
Code-based planning. Another approach to defining robot behaviors using LLM revolves around the direct generation of executable code. An excellent example is Code as Policies [10]. In this work, the authors release a robot-centric formulation of language model generated programs (i.e., LMPs), based on hierarchical code generation prompting, that allows the generation of new policy code via recursively defining undefined functions. LMPs take advantage of few-shot prompting to generate different subprograms and can be generated hierarchically with chain-of-thought prompting. Even if their approach requires no additional training, their system is based on GPT-3 with 175B parameters. Moreover, they use open-vocabulary object detection models like ViLD [5] and MDETR [9] to compose perception-to-control feedback logic. Another work that relies on prompt engineering to create a Pythonic planner is ProgPrompt [16]. They introduce a programmatic LLM prompt structure that enables plan generation functional across situated environments and robot capabilities. Their fine-tuned version of GPT-2 can generate plausible action plans in the context of robotic task planning, this is achieved by providing the model with a Pythonic program-like specification of the the available actions and objects in an environment as well as executable example programs. In this way, they introduce situated awareness in LLM-based robot task planning via prompting.
@@ III Method
For the experiments in this work, we considered as reference models the state-of-the-art open-source large language models provided by Meta. We selected Llama-2-7b [19], the foundation model, Llama-2-7b-chat [19], designed for conversational tasks, and CodeLlama-7b-Instruct [15], with a focus on code-generation. All of these models are based on the transformer architecture pioneered in [20]. Following the approach proposed in [11], we fine-tuned the foundation model Llama-2-7b with the Alpaca dataset, with the procedure reported in [17] using the Alpaca-LoRA repository [21]. For the remainder of this paper, we will refer to this fine-tuned version as Alpaca. Additional fine-tuning of these models allows for customization to meet specific requirements, such as domain specialization, in our case the generation of behavior trees. A summary of the various fine-tuning steps is presented in Figure 1, and a more detailed description is provided in Section III-C.
@@ III-A Dataset Format
Each entry of our instruction-following dataset is composed of three parts: “instruction”, “input” and “output”. While the original work of [17] states that “input” is optional, in our case, we combine the XML version of the behavior tree provided in “output” with the description of the task in the “input” to create a definition that matches the prompts used later during inference.
You will be provided a summary of a task performed by a robot, and your objective is to express this task as a behavior tree in XML format. The behavior tree is a simple sequential task for a robot. It first instructs the robot to move to a specific point (GoPoint) and then to interact with a particular object (GoObject). The robot will execute these tasks in sequence, moving to the specified destination before interacting with the designated object. <root main_tree_to_execute = "MainTree" >   <BehaviorTree ID="MainTree">   <Sequence name="root_sequence">     <GoPoint goal="{destination}"/>     <GoObject target="{object}"/>   </Sequence>   </BehaviorTree> </root>
This is an example of an entry in the dataset. In gray, the “instruction” element is common to all the samples in the dataset, and it is fixed and immutable. It is followed by the “input” element, in blue, where a description in natural language is provided of the behavior tree. Finally, the “output” element represents the behavior tree in XML format. All the behavior trees used in the dataset are sourced via the work of [4]. In this work, the authors provide a collection of roughly 600 behavior trees collected from various open-source projects in the field of robotics. This dataset is particularly valuable because compatible with BehaviorTree.CPP, the de facto standard of ROS2, and the same format used by [11]. Furthermore, what distinguishes this dataset is the quality of the behavior trees, which have been developed for actual applications and tested on robots and thus have already been validated in real-world scenarios. The natural language description (i.e., the “input” element) of each entry of the dataset has been generated automatically using GPT-3.5, with a procedure described in the next section.
@@ III-C Fine-Tuning Process
The models selected for the fine-tuning process are Llama2-7b, LlamaChat, and CodeLlama-Instruct-7b. The fine-tuning process can be structured in two steps, the first one involves the Llama2-7b model that has been fine-tuned with the original Alpaca dataset as reported in [7]. This step prepared the base model for instruction-following tasks, differently from the other models that are already fine-tuned for this task. The second step prepared the models for the generation of the behavior trees. To perform this task, we further fine-tuned the models using the dataset described in III-A. LLMs fine-tuning is a computationally intensive task, therefore, we employed the Parameter-Efficient-Fine-Tuning (PEFT) approach, in particular, Low-Rank Adaptation (LoRA) described in detail in [7]. As reported in [14], this is particularly useful to avoid retraining the entire model from scratch. In this way, we can keep the model frozen and add an adapter at the end of the model consisting of a few learnable parameters, and layers.
Regarding LoRA parameters, most of them were set at their default value, except for the target modules. In the work [7], it is noted that the transformer architecture has four weight matrices in the self-attention module ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$). To enhance generalization capabilities with our limited dataset and to optimize performance, it was necessary to unfreeze the MLP layers. We expanded the corresponding initial set of weight matrices [q_proj, k_proj, v_proj, o_proj] by adding the additional weight matrices of the MLP module, these include [gate_proj, up_proj, down_proj]. This approach led to a more robust and effective training process.
The fine-tuning process was accomplished with two NVIDIA RTX Quadro 6000 with a total of 48GB of video memory. All the basic hyperparameters have been used except for a batch size of 256 and a micro-batch size of 4, we used a learning rate of 3e-4 and a validation set size of 5%. Given that our dataset is composed of less than a thousand samples, we increased the number of epochs to achieve satisfying results.
@@ IV-A Task definition
    Navigation with priority. The system receives a list of locations and a list of corresponding readings (e.g., temperature). Given a threshold, the robot must visit all locations prioritizing those with a reading above the threshold.
@@ IV-B Evaluation approach
The evaluation has been accomplished in two separate phases: the first one represents a preliminary selection among the models while the second one effectively validates the models and the generated behavior trees with more rigorous metrics. In the first phase, we evaluated the models on a subset of simpler tasks, namely tasks from 1 to 7. In the second phase, where the general capabilities of the models are already been assessed, all nine tasks are considered. The metrics considered in the first evaluation phase are:
@@
  • Semantic Evaluation: without considering the syntactic correctness, we evaluate the semantic accuracy. This means the ability of the LLM to generate a solution that understands and solves the problem at hand. This evaluation is performed by a group of human experts.
Figure 2: Simulation Environment using TurtleBot3 on Gazebo
@@
  • Simulation robot: verifying the successful execution of the task on a simulated environment, in particular, the navigation task has been tested on the TurtleBot3 by ROBOTIS. See Figure 2.
@@
TABLE I: Time Comparison
Alpaca LlamaChat CodeLlama
Base, Zero-Shot
Mean - 3m 10s 2m 52s
Std. Dev. - 1m 29s 1m 57s
Fine Tuned, Zero-Shot
Mean - 1m 13s 1m 15s
Std. Dev. - 56s 58s
Base, One-Shot
Mean 1m 42s 3m 54s 1m 49s
Std. Dev. 46s 38s 1m 16s
Fine Tuned, One-Shot
Mean 2m 10s 1m 25s 1m 21s
Std. Dev. 1m 20s 58s 1m 11s
@@
IV-C Preliminary selection
Table I shows the time comparison of execution time while doing inference on the model, the tests have been conducted on an NVIDIA RTX Quadro 6000. Considering the models without fine-tuning. When using a zero-shot approach, CodeLlama overall is faster than LlamaChat, having a similar mean but less standard deviation. In general, Alpaca is slower and task 2 and task 7 do not terminate, therefore, it was not possible to compute the mean execution time. One-shot prompts are slightly slower since they have to deal with a more complex input, CodeLlama and LlamaChat have comparable mean and standard deviation similar to their zero-shot counterparts. It is worth noting that LlamaChat becomes rather consistent (i.e., low standard deviation) between tasks when using a one-shot prompt. In this case, Alpaca receives a significant performance boost, since it can complete all tasks, however it remains slower in most cases.
Regarding syntactic correctness, we define a behavior tree to be syntactically correct if successfully recognized by Groot2. Groot2 checks the correctness of the overall tree for what concerns Nodes, Decorators, and parameters. All the outputs of three base models and the corresponding fine-tuned version with our dataset, both with zero-shot and one-shot prompting, have been evaluated using Groot2. The results of this evaluation are summarised in Table II.
In terms of accuracy, base models with a zero-shot prompt achieve very limited results, with only CodeLlama being able to generate a syntactically correct behavior tree in a very limited number of cases. Adding examples in the prompt (i.e., one-shot) significantly boosts performance, with CodeLlama managing to create a syntactically correct behavior tree for each task. A similar result is achieved for fine-tuned models. In this case, even with a zero-shot prompt, all models can generate syntactically correct behavior trees for most tasks. Surprisingly, providing an example does not drastically increase accuracy as before, with the extreme case of CodeLlama achieving a worse result than before. This result shows how impactful the choice of examples can be in the generation of the final result.

### v2

@@ I Introduction
An example of these representations is behavior trees (BTs). Behavior trees have gained significant traction in robotics, offering a structured and scalable approach to managing complex robot behaviors. Originating from the video game industry, BTs have found relevance in robotics due to their ability to handle high-level decision-making and control in dynamic environments. A key player in this transition is the BehaviorTree.CPP library [^1] https://www.behaviortree.dev/. This library has become the de facto standard in the Robot Operating System 2 (ROS2) ecosystem, largely due to its inclusion in Navigation2 [12]. This integration has facilitated the development of sophisticated navigation behaviors, contributing to the broader adoption and standardization of BTs in robotics software design.
@@
Recently, large language models (LLM) trained on large amounts of data have emerged as a promising tool to manage this challenge and deal with the complex knowledge required for task planning. These models have demonstrated remarkable success in robot planning from natural language instructions. Additionally, an LLM can represent commonsense priors (e.g., visit the colder locations first) and comprehend spatial relationships (e.g., move behind the sofa). However, such large models present a major limitation: their substantial hardware resource requirements prevent their direct deployment on robots.
This work explores an alternative approach, focusing on compact large language models, specifically those with 7 billion parameters. We posit that these smaller and less resource-intensive models can generate behavior trees for robots with comparable effectiveness to their larger counterparts. Our main contributions are: (i) a dataset of 600 behavior trees paired with a natural language description of their tasks, (ii) a system to evaluate and validate the performance of the generated BT over nine different tasks including navigation and manipulation, and (iii) a comparative analysis of three different LLM and their fine-tuned versions. All the material used in this work, including the dataset, the complete prompts of both evaluation phases, and the source code of the validator are available at https://github.com/AIRLab-POLIMI/BTGenBot.
@@ II Related work
A pre-trained foundation model (PFM) is a complex neural network-based model trained on extensive open-source datasets, boasting billions of parameters [23]. This robust architecture serves as a versatile backbone that can be further fine-tuned for diverse tasks across various domains. Recently, several foundation models specifically trained on massive text datasets (i.e., large language models) have been proposed such as the most recent one GPT-4 [1] by OpenAI or LLaMA by Meta AI [18]. In this work, we rely on LLaMA released by Meta AI, a family of foundation models trained on an open-source dataset with trillions of tokens. In particular, we employed LLaMA-2 [19], trained on 2 trillion tokens, and with a context length of 4000 tokens, double the length of its predecessor. Three model sizes are available: 7, 13, and 70 billion parameters.
@@
Behavior tree generation. An example of behavior tree generation using LLM is the work done in [11]. In this paper, the authors exploit a transformer-based LLM fine-tuned from the Stanford Alpaca 7B model [17] to generate behavior trees from a text description. They further fine-tuned this model with natural language descriptions generated using text-davinci-003 from behavior trees created in the previous step. In summary, they use two different LLMs to generate both the behavior trees and the task description. While interesting in proving that LLMs can be used to generate BTs, this approach has several limitations. Fine-tuning an LLM with a dataset generated by the LLM itself causes a data bias that is only partially mitigated by using an auxiliary LLM to generate the descriptions. Moreover, self-generated data can be noisy and, in this configuration, errors propagate from the data generation phase to the training phase. Differently from the solution proposed in [11], we create our instruction-following dataset by collecting behavior trees from open-source robotics projects. This allows us to fine-tune our model with BTs already tested and effectively used in real projects. Moreover, we use GPT-3.5 only to generate descriptions for each corresponding behavior tree.
@@
Finally, concerning validating the above approach, the only metric used is the visual evaluation by a group of human experts. The evaluators were presented with a pair of behavior trees and had to indicate which one was generated. This evaluation, while interesting and useful to asses the general capabilities of the model, is rather limited. This method does not consider the correctness from a syntactical and semantic point of view, or any performance metric, such as the generation time and the hardware resources used. In contrast, our approach aims at creating fully functional behavior trees to be used in real-world applications. Therefore, we tested the generated outputs with the metrics described in Section III.
@@
Code-based planning. Another approach to defining robot behaviors using LLM is to generate executable code. An excellent example is Code as Policies [10]. In this work, the authors release a robot-centric formulation of language model generated programs (i.e., LMPs), based on hierarchical code generation prompting, that allows the generation of new policy code via recursively defining undefined functions. LMPs use few-shot prompting to generate different subprograms and can be generated hierarchically with chain-of-thought prompting. Even if their approach requires no additional training, their system is based on GPT-3 with 175B parameters. Moreover, they use open-vocabulary object detection models like ViLD [5] and MDETR [9] to compose perception-to-control feedback logic. Another work that relies on prompt engineering to create a Pythonic planner is ProgPrompt [16]. They introduce a programmatic LLM prompt structure that enables plan generation functional across situated environments and robot capabilities. Their fine-tuned version of GPT-2 can generate plausible action plans in the context of robotic task planning, this is achieved by providing the model with a Pythonic program-like specification of the the available actions and objects in an environment and executable example programs. In this way, they introduce situated awareness in LLM-based robot task planning via prompting.
@@ III Method
For the experiments in this work, we considered the state-of-the-art open-source large language models provided by Meta as reference models. We selected Llama-2-7b [19], the foundation model, Llama-2-7b-chat [19], designed for conversational tasks, and CodeLlama-7b-Instruct [15], with a focus on code-generation. These models are based on the transformer architecture pioneered in [20]. Following the approach proposed in [11], we fine-tuned the foundation model Llama-2-7b with the Alpaca dataset, with the procedure reported in [17] using the Alpaca-LoRA repository [21]. For the remainder of this paper, we will refer to this fine-tuned version as Alpaca. Additional fine-tuning of these models allows for customization to meet specific requirements, such as domain specialization (e.g., the generation of behavior trees). A summary of the various fine-tuning steps is presented in Figure 1, and a more detailed description is provided in Section III-C.
@@ III-A Dataset Format
Each entry of our instruction-following dataset has three parts: “instruction”, “input” and “output”. While the original work of [17] states that “input” is optional, in our case, we combine the XML version of the behavior tree provided in “output” with the description of the task in the “input” to create a definition that matches the prompts used later during inference.
{mdframed}
[linecolor=light-gray] You will be provided a summary of a task performed by a robot, and your objective is to express this task as a behavior tree in XML format.
The behavior tree is a simple sequential task for a robot. It first instructs the robot to move to a specific point (GoPoint) and then to interact with a particular object (GoObject). The robot will execute these tasks in sequence, moving to the specified destination before interacting with the designated object. {mdframed}[backgroundcolor=highlight-example, innertopmargin=-4pt, innerbottommargin=-4pt, innerleftmargin=2pt, innerrightmargin=2pt]
<root main_tree_to_execute = "MainTree" >
<BehaviorTree ID="MainTree">
<Sequence name="root_sequence">
<GoPoint goal="{destination}"/>
<GoObject target="{object}"/>
</Sequence>
</BehaviorTree>
</root>
This is an example of an entry in the dataset. In gray, the “instruction” element is common to all the samples in the dataset, and it is fixed and immutable. It is followed by the “input” element, in blue, where a description in natural language is provided of the behavior tree. Finally, the “output” element represents the behavior tree in XML format. All the behavior trees used in the dataset are sourced via the work of [4]. In this work, the authors provide a collection of roughly 600 behavior trees collected from various open-source projects in the field of robotics. This dataset is particularly valuable because compatible with BehaviorTree.CPP, the de facto standard of ROS2, and the same format used by [11]. Furthermore, what distinguishes this dataset is the quality of the behavior trees, which have been developed for actual applications and tested on robots and thus have already been validated in real-world scenarios. The natural language description (i.e., the “input” element) of each dataset entry has been generated automatically using GPT-3.5, with a procedure described in the next section.
@@ III-C Fine-Tuning Process
The models selected for the fine-tuning process are Llama2-7b, LlamaChat, and CodeLlama-Instruct-7b. The fine-tuning process can be structured in two steps, the first involves the Llama2-7b model that has been fine-tuned with the original Alpaca dataset as reported in [7]. This step prepared the base model for instruction-following tasks, differently from the other models already fine-tuned for this task. The second step prepared the models for the generation of the behavior trees. We further fine-tuned the models to perform this task using the dataset described in III-A. LLMs fine-tuning is a computationally intensive task, therefore, we employed the Parameter-Efficient-Fine-Tuning (PEFT) approach, in particular, Low-Rank Adaptation (LoRA) described in detail in [7]. As reported in [14], this is particularly useful to avoid retraining the entire model from scratch. This way, we can keep the model frozen and add an adapter at the end of the model consisting of a few learnable parameters, and layers.
Most LoRA parameters were set at their default value, except for the target modules. In the work [7], it is noted that the transformer architecture has four weight matrices in the self-attention module ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$). To enhance generalization capabilities with our limited dataset and to optimize performance, it was necessary to unfreeze the MLP layers. We expanded the corresponding initial set of weight matrices [q_proj, k_proj, v_proj, o_proj] by adding the additional weight matrices of the MLP module, these include [gate_proj, up_proj, down_proj]. This approach led to a more robust and effective training process.
The fine-tuning process was accomplished with two NVIDIA RTX Quadro 6000 with a total of 48GB of video memory. All the basic hyperparameters have been used except for a batch size of 256 and a micro-batch size of 4, we used a learning rate of 3e-4 and a validation set size of 5%. Given that our dataset contains less than a thousand samples, we increased the number of epochs to achieve satisfying results.
@@ IV-A Task definition
    Navigation with priority. The system receives a list of locations and a list of corresponding readings (e.g., temperature). The robot must visit all locations prioritizing those with a reading above a given threshold.
@@ IV-B Evaluation approach
The evaluation has been accomplished in two phases: first a preliminary selection among the models, and then a validation of the models and the generated behavior trees with more rigorous metrics. In the first phase, we evaluated the models on a subset of simpler tasks (i.e., tasks 1 to 7). In the second phase, where the general capabilities of the models are already been assessed, all nine tasks are considered. The metrics considered in the first evaluation phase are:
@@
  • Semantic Evaluation: without considering the syntactic correctness, we evaluate the semantic accuracy. This means the ability of the LLM to generate a solution that understands and solves the problem at hand. A group of human experts performs this evaluation.
@@
  • Simulation robot: verifying the successful execution of the task on a simulated environment, in particular, the navigation task has been tested on the TurtleBot3.
@@
TABLE I: Performance reference
GPT-4o Gemini Llama2 13B
Task 1
Task 2
Task 3
Task 4
Task 5
Task 6
Task 7
Task 8
Task 9
@@
IV-C Performance reference
To provide a frame of reference for the performance achieved by our approach, we tested our nine tasks against state-of-the-art LLMs. We selected GPT-4o, Gemini, and Llama 2 13B Chat. Table I shows the results. The values reported in the table are obtained using the same one-shot prompts used in our evaluations. The generated behavior trees are validated with our custom validator without any post-processing. GPT-4o achieves the best performance, it can generate a correct solution for seven out of nine tasks. Gemini can provide a working solution only in four tasks. While, as expected, behavior trees generated by Llama 2 13B Chat are correct only in tasks 1 and 4, which are the less challenging. This simple evaluation shows that even larger LLMs may struggle to generate syntactically and semantically correct behavior trees without an accurate prompt or fine-tuning. In particular, in some cases, it is enough to fix small syntactical errors (e.g., correct a Decorator name) to achieve a correct solution.
IV-D Preliminary selection
Since this approach is intended to be used directly on the robot, we compare the inference times of each model when running on a Jetson AGX Orin 64GB. The inference time ranges from a few seconds for simpler tasks to a few minutes when considering more complex prompts. This pattern is confirmed between zero-shot and one-shot prompts, where the longer prompt (i.e., one-shot) has a longer inference time. When discussing inference time, it is worth noting that in some cases, Alpaca fails to terminate the generation process.
Regarding syntactic correctness, we define a behavior tree as syntactically correct if successfully recognized by Groot2. All the outputs of three base models and the corresponding fine-tuned version with our dataset, both with zero-shot and one-shot prompting, have been evaluated using Groot2. The results of this evaluation are summarised in Table II. Base models with a zero-shot prompt achieve poor results, only CodeLlama can generate a syntactically correct behavior tree in a few cases. Adding examples in the prompt (i.e., one-shot) significantly boosts performance, with CodeLlama creating a syntactically correct behavior tree for each task. A similar result is achieved for fine-tuned models. In this case, even with a zero-shot prompt, all models can generate syntactically correct behavior trees for most tasks. Surprisingly, providing an example does not drastically increase accuracy as before, with the extreme case of CodeLlama achieving a worse result than before. This result shows how impactful the choice of examples can be on the final output.